# Robust Signal Detection with Quadratically Convex Orthosymmetric Constraints

### v2

@@ Abstract
This paper is concerned with robust signal detection in Gaussian noise under quadratically convex orthosymmetric (QCO) constraints. Specifically, the null hypothesis $H_{0}$ assumes no signal, whereas the alternative $H_{1}$ considers signal that is separated in Euclidean norm from zero and belongs to a set $K$ satisfying the QCO constraints. In addition, an adversary is allowed to inspect all the $N$ samples and replace up to $\epsilon N$ of them with arbitrary values, where $0<\epsilon<c_{0}<\frac{1}{2}$ is the corruption rate. Our main results establish the minimax rate of the separation radius $\rho_{\text{critical}}$ between $H_{0}$ and $H_{1}$ purely in terms of the geometry of $K$, the corruption rate $\epsilon$ (up to logarithmic factors in $\frac{1}{\epsilon}$) and the scale of the noise $\sigma$. We argue that the Kolmogorov widths of the constraint play a central role in determining the minimax rate. This indicates similarity with the (uncorrupted) estimation problem under QCO constraints, which was first established by Donoho et al., (1990). Moreover, the minimax lower bound reveals interesting phase transitions of the testing problem regarding $\epsilon$. Consistent with classic belief about testing and estimation, the testing problem is “easier” (even when one compares the results with recent papers studying the constrained robust estimation problem (Prasadan and Neykov,, 2025)). In addition to the main results above, where the upper bound is achieved with an intractable algorithm, inspired by Canonne et al., (2023), we develop a polynomial time algorithm which also nearly (up to logarithmic factors in $N$ and $1/\epsilon$) achieves the minimax lower bound. In contrast to Canonne et al., (2023), our algorithm works for signals of arbitrary Euclidean length, and respects the QCO constraint. Finally, all the results above are naturally extended to the $\ell_{p}$ norm testing problem for $1\leq p<2$.
@@ 1 Introduction
Performing inference under constraints and even contamination could be a formidable task. Most of the related works focus on the estimation and testing problem under several simple constraints with no contamination (Donoho and Gasko, (1992), Liu and Moitra, (2021), Liu and Moitra, (2023)). One crucial reason for this is that most of the methods and techniques heavily depend on the assumption of the true model behind the data (for example, likelihood ratio test and tests depending on the asymptotic distributions). Consider the following special case where $K=\mathbb{R}^{d}$ and $\epsilon=0$. Given $N$ samples $X_{1},X_{2},\dots,X_{N}\in\mathbb{R}^{d}$ from $\mathcal{N}(\mu,\sigma^{2}\mathbf{I}_{d})$ with unknown $\mu$ and known $\sigma\in\mathbb{R}^{+}$. In the testing problem, we are asked to distinguish between two hypotheses:
@@
In the estimation problem, we are asked to give an estimate $\widehat{\mu}$ of $\mu$ to achieve a small error under a certain loss function (a natural one here is the $\ell_{2}$ loss). $\bar{X}=\frac{1}{N}\sum\limits_{i=1}^{N}X_{i}$ is a sufficient statistic for the model. The testing and estimation based on $\bar{X}$ is widely known to be optimal for both problems. Specifically, to control errors within the level of (small) constant, one needs $\rho\gtrsim\frac{d^{1/4}}{\sqrt{N}}\sigma$ for the testing problem (in the minimax sense) (Ingster and Suslina, (2003), Diakonikolas et al., (2022)) while $\inf\limits_{\widehat{\mu}}\sup\limits_{\mu\in\mathbb{R}^{d}}\mathbb{E}\left\lVert\widehat{\mu}-\mu\right\rVert_{2}^{2}\gtrsim\frac{d}{N}\sigma^{2}$ if one wants to optimally estimate $\mu$ in $\ell_{2}$ norm (LeCam, (1973), also see Example 15.4 in Wainwright, (2019)). However, as mentioned, the property of $\bar{X}$ is very sensitive to the authenticity of the samples. The method quickly breaks down with even modification on a single sample (the adversary may modify any single datum, thereby forcing the mean to be zero and causing considerable type-II error.). This is precisely the major reason why testing and estimation problems are very challenging when studied from a robust perspective.
@@
Introduced by Kolmogoroff, (1936) in the study of function spaces and approximation theory, the Kolmogorov widths (also known as the Kolmogorov $k$-width) characterizes how well an element can be reached by some linear approximations of intrinsic dimension $k$, given that the element belongs to a linear space and some restrictions are applied. For a given $k$, it selects the best $k$-dimensional projection operator under the measure of minimax error. The reader can refer to Definition 2.2 for a formal presentation. Kolmogorov $k$-width is a classic concept and has seen usage in various fields. See Floater et al., (2021) for a recent work in the problem of low-rank approximation of matrices using the Kolmogorov $k$-width. Evans et al., (2009) uses this concept to study the $k$-method in isogeometric analysis. Kolmogorov $k$-width is also popular in the problems related to partial differential equations (PDEs), where the decay rate of the Kolmogorov $k$-width is intensely studied (Cohen and DeVore, (2015), Greif and Urban, (2019)). See also Papapicco et al., (2022) for an application in neural networks and PDEs, and a very recent work, Florian Arbes and Urban, (2025), for an application in optimal transportation and PDEs. Finally, see Greif and Urban, (2019) and Liang and Parks, (1996) for its power in the wave problem and wavelet representations. One can find more details and properties of the Kolmogorov $k$-width in Pinkus, (2012).
@@ 1.1 Problem Formulation and Our Contributions
XX0KK
Figure 1: Testing $H_{0}:\mu=0$ vs $H_{1}:\|\mu\|_{2}\geq\rho,\mu\in K$ for the axis aligned ellipsoid (which is a special case of a QCO set) $K$ based on a single observation $X$. Upon applying a rotation to the data, one can use the test we develop in Section 3.2 to solve the problems optimally for rotated (non-axis aligned) ellipsoids or more generally QCO sets (which are still centered around the origin). See our first remark in the discussion section for more details.
Our main result shows that the $\rho_{\text{critical}}$ is determined by three independently valid lower bounds. Moreover, the lower bounds are tight and optimal with only the exception of logarithmic terms. The first lower bound captures the geometric nature of the QCO set $K$ via the Kolmogorov $k$-width. This lower bound is independent of the adversary, and also works in the uncorrupted setting, i.e., when $\epsilon=0$. The second lower bound is obtained within the framework of differential privacy, which is introduced in Section B.3. The third lower bound is derived from a combined argument related to identical mixtures, computation of the total variation and a designed strategy of the adversary. For the upper bounds, we present two algorithms which conclude whether $H_{0}$ holds or not while ensuring type-I and type-II errors both asymptotically less than $\alpha$. The requirements on $\lVert\mu\rVert_{2}$ imposed by the first algorithm are slightly closer to the theoretical lower bounds; however, this improvement comes at the cost of exponential time complexity, rendering the method computationally intractable for large $N$ and $d$. The second algorithm, in contrast, provides a slightly looser upper bound than the first algorithm (though still nearly optimal) while achieving polynomial time complexity. Nevertheless, both upper bounds suggest that the maximum of the three lower bounds discussed previously suffices to fully characterize the behavior of $\rho_{\text{critical}}$. Finally, the results and algorithms derived in the $\ell_{2}$ norm scenario can be naturally extended to the $\ell_{p}$ case for $1\leq p<2$. Section 5 contains such relevant discussions.
@@ 1.2 Related Literature
In this section, we discuss some relevant works. We first comment on the literature on minimax testing with uncorrupted data. In their seminal body of work on nonparametric testing, Ingster, (1982); Ingster, 1993a ; Ingster, 1993b ; Ingster, 1993c ; Ingster and Suslina, (2000) showed the difference between testing and estimation in multiple problems. See also Ingster and Suslina, (2003) for a summary of the many existing results in the area. Despite various progress in goodness-of-fit testing in Gaussian models, we are not aware of a general result dedicated to QCO sets. For certain special types of ellipsoids, the problem was first studied by Ermakov, (1991), where the author derived exact minimax rates. Suslina, (1996) studied the asymptotics of the problem of signal detection in $\ell_{q}$ ellipsoids with $\ell_{p}$ ball removed. Since $\ell_{q}$ ellispoids for $q\geq 2$ are QCO sets, when $p=2$ we provide a generalization of these results. For general ellipsoids, Baraud, (2002) studied the testing problem in $\mathbb{R}^{n}$ and $\ell_{2}$ and obtained the minimax optimal critical radius. Our work can hence be seen as a generalization to part of Baraud, (2002). In an interesting recent work, Wei and Wainwright, (2020) studied local testing rates on ellipsoids. Here local means that the point of interest under the null is not necessarily the zero point, but can be an arbitrary point in the ellipsoid. Wei and Wainwright, (2020) proved upper and lower bounds based on localized Kolmogorov widths, which are not guaranteed to match in general, but do match for certain examples with Sobolev ellipsoids. In contrast, the present work focuses on more general sets, but the point of interest under the null hypothesis is always anchored at the zero point, and the critical radius is determined based on the global (instead of local) Kolmogorov widths of the set.
@@ 1.3 Organization
This paper is structured as follows. Section 2 introduces the notations and several important concepts used through the paper. Section 3 presents the lower and upper bounds for the signal detection problem and contains the main results of the paper. It also contains the discussion of the phase transitions of the problem. Section 4 presents the experimental results of the polynomial time algorithms related to Theorem 3.8 in Section 3. Section 5 contains the discussion which compares the minimax rates of testing to estimation and the extension of the main results to the $\ell_{p}$ norm separation. It also includes some open problems for future works. We attach necessary probabilistic tools and other lemmas in Appendix A. The proofs to the lower and upper bounds and other main corollaries in Section 3 and Section 5 are included in Appendix B. Appendix C presents the pseudo codes of related algorithms.
@@ 2.1 Notations and Definitions
To simplify the analysis, we assume that $K\subset\mathcal{X}=\mathbb{R}^{d}$, and therefore $\mu\in\mathbb{R}^{d},X_{i}\in\mathbb{R}^{d}$ through the paper. However, one can extend the results to the case when $K\subset\ell_{2}$ with the same techniques and algorithms with minimal efforts given the assumption $\lim\limits_{k\rightarrow\infty}D_{k}(K)=0$. See Definition 2.2 for further details. We use $\rho$ to denote the value of the $\ell_{2}$ norm of $\mu$ under $H_{1}$. $\left\lVert\cdot\right\rVert_{p}$ means the $\ell_{p}$ norm in $\mathbb{R}^{d}$ for $1\leq p\leq\infty$. When $A$ is a set, $\left|A\right|$ denotes the cardinality of $A$. We use calligraphic font $\mathcal{P},\mathcal{Q},\mathcal{R}\dots$ to represent probability distributions. $\mathcal{P}\ll\mathcal{Q}$ means that $\mathcal{P}$ is absolutely continuous with respect to $\mathcal{Q}$. $\text{TV}(\mathcal{P},\mathcal{Q})$ denotes the total variation distance and $\text{KL}(\mathcal{P},\mathcal{Q})$ for the Kullback-Leibler divergence if $\mathcal{P}\ll\mathcal{Q}$. When the underlying probability distribution is not explicitly specified, we use $\mathbb{P}$ alternatively to denote probability law in general. We use $\left\langle\cdot,\cdot\right\rangle$ and $(\cdot)^{\top}(\cdot)$ alternatively (possibly with a subscript when needed) to represent the inner product in a Hilbert space. $\text{diag}\{a\}$ represents the diagonal matrix with diagonal entries $a_{i},i=1,2,\dots$. The convenient notation $[n]$ denotes the set $\left\{1,2,\dots,n\right\}$. When $I$ is an index set, $v_{I}$ represents the restriction of $v$ on $I$ if $v$ is a vector and $\mathbf{Y}_{I}$ represents the sub-matrix of $\mathbf{Y}$ with rows selected according to $I$ if $\mathbf{Y}$ is a matrix. $\mathcal{O}$, $\varTheta$, $\varOmega$ notations are used to denote the asymptotic order of rates. The lowercase $c_{1},c_{2},\dots$ denote the universal constants and their specific values may differ from line to line.
@@
Given a set $K\subset\mathbb{R}^{d}$, we say it is a quadratically convex orthosymmetric (QCO) set if it satisfies the following conditions:
 (1). $K$ is convex;
 (2). $K$ is quadratically convex, which means that $K^{2}$ is also convex, where $K^{2}$ is defined as
@@
(3). $K$ is orthosymmetric, which means that if $\theta=\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K$, then $\theta_{\eta}:=\left(\eta_{1}\theta_{1},\dots,\eta_{d}\theta_{d}\right)^{\top}\\ \in K$, where $\eta_{i}\in\left\{-1,1\right\},1\leq i\leq d$.
@@
The definition of QCO set can be naturally extended to the case when $K\subset\ell_{2}$.
Definition 2.2 (Kolmogorov $k$-width).
@@
In this work, our analysis will be confined to orthogonal projections. Moreover, though $\mathcal{P}_{k}$ is the set of all orthogonal projection with intrinsic dimension $k$, we can discretize this set — only consider the projections that are aligned with the coordinate axes without loss of generality, as shown by the analysis later. For now, let us first assume that the norm leveraged in the definition is the $\ell_{2}$ norm.
Based on the definition, one can derive several basic properties of the Kolmogorov widths. It is non-increasing with respect to $k$, meaning that $D_{k_{1}}(K)\geq D_{k_{2}}(K)$ for any $0\leq k_{1}\leq k_{2}$. In addition, for a finite-dimensional space $\mathcal{X}$ with full dimension $d$, we always have $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}$ and $D_{d}(K)=0$ for any $K\subset\mathcal{X}$. For the more general case where $K\subset\ell_{2}$, we shall assume $\lim\limits_{d\rightarrow\infty}D_{d}(K)=0$. In this paper, we further define $D_{k}(K)=\infty$ when $k<0$ for the convenience of the marginal cases.
The connection between the Kolmogorov widths and our main results is established through the following definition of the first and second dimensions of the optimal projections, which are designed to identify the optimal $k$ satisfying certain marginal conditions determined by the geometric properties of the constraint set and the corruption process.
@@
$H_{d}(a_{1},\dots,a_{d}):=\left\{x\left\lvert\right.x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d},\left|x_{i}\right|\leq a_{i},a_{i}>0,1\leq i\leq d\right\}.$
Obviously, $H_{d}(a_{1},\dots,a_{d})$ is convex and orthosymmetric. One can verify that $H_{d}^{2}(a_{1},\dots,a_{d})=\left\{x^{2}\left\lvert\right.x\in H_{d}(a_{1},\dots,a_{d})\right\}$ is also a hyperrectangle with each side being $0,a_{i}^{2}],1\leq i\leq d$. Therefore, it is a QCO set. By reindexing if necessary, we may assume that the sequence $\left\{a_{i}\right\}_{i=1}^{d}$ is arranged in decreasing order. The Kolmogorov $k$-width is
@@
Similarly, we assume $\left\{a_{i}\right\}$ is arranged in decreasing order, by which the Kolmogorov widths are
@@
The first optimal dimension is $k_{1}^{\star}=\max\limits\left\{j\left\lvert\right.1\leq j\leq d,a_{j}>\frac{\sqrt{j}}{N}\sigma^{2}\right\}$ when $a_{1}>\frac{1}{N}\sigma^{2}$ and 0 otherwise. The second optimal dimension $k_{2}^{\star}=\max\limits\left\{j\left\lvert\right.1\leq j\leq d,a_{j}>\frac{\sqrt{j}\epsilon}{\sqrt{N}}\sigma^{2}\right\}$ when $a_{1}>\frac{\epsilon}{\sqrt{N}}\sigma^{2}$ otherwise 0.
@@
When a ellipse is not axes-aligned, (1), the orthosymmetric property is not guaranteed, which can be easily checked; (2), the quadratically convex property is true for $d=2$ and false for $d\geq 3$. To gain the idea of (2), consider an ellipse

### v3

@@ Abstract
This paper studies the problem of robust signal detection in Gaussian noise under quadratically convex orthosymmetric (QCO) constraints. We consider a minimax testing framework where the signal belongs to a QCO set and is separated from zero in Euclidean norm, while an adversary is allowed to arbitrarily corrupt a fraction $\epsilon$ of the samples. We establish the minimax separation radius between the null and alternative purely in terms of the constraint geometry, sample size, corruption rate, and noise scale. Our analysis argues that the Kolmogorov widths of the constraint set play a central role in determining the detection limits, paralleling to classic results in estimation problem. The derived lower bounds exhibit phase transitions with respect to the corruption rate and confirm that robust testing is statistically easier than robust estimation. While the information-theoretic upper bound is achieved by a computationally intractable test, we develop a polynomial-time algorithm that achieves the minimax lower bound up to logarithmic factors. Unlike prior work, our algorithm handles signals of arbitrary Euclidean length while respecting the QCO constraints. Finally, we extend these results to the robust $\ell_{p}$ norm testing for $1\leq p<2$.
@@ 1 Introduction
Performing inference under constraints and even contamination could be a formidable task. Most of the related works focus on the estimation and testing problem under several simple constraints with no contamination (Donoho and Gasko, (1992); Liu and Moitra, (2021, 2023)). One crucial reason for this is that most of the methods and techniques heavily depend on the assumption of the true model behind the data (for example, likelihood ratio test and tests depending on the asymptotic distributions). Consider the following special case where $K=\mathbb{R}^{d}$ and $\epsilon=0$. Given $N$ samples $X_{1},X_{2},\dots,X_{N}\in\mathbb{R}^{d}$ from $\mathcal{N}(\mu,\sigma^{2}\mathbf{I}_{d})$ with unknown $\mu$ and known $\sigma\in\mathbb{R}^{+}$. In the testing problem, we are asked to distinguish between two hypotheses:
@@
In the estimation problem, we are asked to give an estimate $\widehat{\mu}$ of $\mu$ to achieve a small error under a certain loss function (a natural one here is the $\ell_{2}$ loss). $\bar{X}=\frac{1}{N}\sum\limits_{i=1}^{N}X_{i}$ is a sufficient statistic for the model. The testing and estimation based on $\bar{X}$ is widely known to be optimal for both problems. Specifically, to control errors within the level of (small) constant, one needs $\rho\gtrsim\frac{d^{1/4}}{\sqrt{N}}\sigma$ for the testing problem (in the minimax sense) (Ingster and Suslina, (2003); Diakonikolas et al., (2022)) while $\inf\limits_{\widehat{\mu}}\sup\limits_{\mu\in\mathbb{R}^{d}}\mathbb{E}\left\lVert\widehat{\mu}-\mu\right\rVert_{2}^{2}\gtrsim\frac{d}{N}\sigma^{2}$ if one wants to optimally estimate $\mu$ in $\ell_{2}$ norm (LeCam, (1973), also see Example 15.4 in Wainwright, (2019)). However, as mentioned, the property of $\bar{X}$ is very sensitive to the authenticity of the samples. The method quickly breaks down with even modification on a single sample (the adversary may modify any single datum, thereby forcing the mean to be zero and causing considerable type-II error.). This is precisely the major reason why testing and estimation problems are very challenging when studied from a robust perspective.
@@
Introduced by Kolmogoroff, (1936) in the study of function spaces and approximation theory, the Kolmogorov $k$-width (also known as the Kolmogorov $N$-width) characterizes how well an element can be reached by some linear approximations of intrinsic dimension $k$, given that the element belongs to a linear space and some restrictions are applied. For a given $k$, it selects the best $k$-dimensional projection operator under the measure of minimax error. Readers can refer to Definition 2.2 for a formal presentation. Kolmogorov $k$-width is a classic concept and has seen usage in various fields. See Floater et al., (2021) for a recent work in the problem of low-rank approximation of matrices using the Kolmogorov $k$-width. Evans et al., (2009) uses this concept to study the $k$-method in isogeometric analysis. Kolmogorov $k$-width is also popular in the problems related to partial differential equations (PDEs), where the decay rate of the Kolmogorov $k$-width is intensely studied (Cohen and DeVore, (2015); Greif and Urban, (2019)). See also Papapicco et al., (2022) for an application in neural networks and PDEs, and a very recent work, Florian Arbes and Urban, (2025), for an application in optimal transportation and PDEs. Finally, see Greif and Urban, (2019) and Liang and Parks, (1996) for its power in the wave problem and wavelet representations. One can find more details and properties of the Kolmogorov $k$-width in Pinkus, (2012).
@@ 1.1 Problem Formulation and Our Contributions
Figure 1: Testing $H_{0}:\mu=0$ vs $H_{1}:\|\mu\|_{2}\geq\rho,\mu\in K$ for the axis aligned ellipsoid (which is a special case of the QCO set) $K$ based on a single observation $X$. Upon applying a rotation to the data, one can use the test we develop in Section 3.2 to solve the problems optimally for rotated (non-axis aligned) ellipsoids or more generally QCO sets (which are still centered around the origin). See our first remark in the discussion section for more details.
Our main result shows that $\rho_{\text{critical}}$ is determined by three independently valid lower bounds. Moreover, the lower bounds are tight and optimal with only the exception of logarithmic terms. The first lower bound captures the geometric nature of the QCO set $K$ via the Kolmogorov $k$-width. This lower bound is independent of the adversary, therefore also works in the uncorrupted setting, i.e., when $\epsilon=0$. The second lower bound is obtained within the framework of differential privacy, which is introduced in Section 3.1.2. The third lower bound is derived from a combined argument related to the identical mixture distributions, computation of the total variation and a designed strategy of the adversary. For the upper bounds, we present two algorithms which conclude whether $H_{0}$ holds or not while ensuring type-I and type-II errors both less than $\alpha$. The requirements on $\lVert\mu\rVert_{2}$ imposed by the theoretical algorithm are slightly closer to the lower bounds. However, this improvement comes at the cost of exponential time complexity, rendering the method computationally intractable for large $N$ and $d$. The polynomial-time algorithm, in contrast, provides a slightly weaker upper bound than the theoretical algorithm (though still nearly optimal) while achieving polynomial-time complexity. Nevertheless, both upper bounds suggest that the three lower bounds previously discussed suffice to fully characterize the behavior of $\rho_{\text{critical}}$. Moreover, such minimax optimal rates also imply a phase transition phenomenon regarding $\epsilon$ for the contamination process. Finally, the results and algorithms derived in the $\ell_{2}$ norm scenario can be naturally extended to the $\ell_{p}$ case for $1\leq p<2$. Section 6 contains such relevant discussions.
@@ 1.2 Related Literature
In this section, we discuss some relevant works. We first comment on the literature on minimax testing with uncorrupted data. In their seminal body of work on nonparametric testing, Ingster, (1982); Ingster, 1993a ; Ingster, 1993b ; Ingster, 1993c ; Ingster and Suslina, (2000) showed the difference between testing and estimation in multiple problems. See also Ingster and Suslina, (2003) for a summary of the many existing results in the area. Despite various progress in goodness-of-fit testing in Gaussian models, we are not aware of a general result dedicated to QCO sets. For certain special types of ellipsoids, the problem was first studied by Ermakov, (1991), where the author derived exact minimax rates. Suslina, (1996) studied the asymptotics of the problem of signal detection in $\ell_{q}$ ellipsoids with $\ell_{p}$ ball removed. Since $\ell_{q}$ ellipsoid for $q\geq 2$ are QCO sets, when $p=2$ we provide a generalization of these results. For general ellipsoids, Baraud, (2002) studied the testing problem in $\mathbb{R}^{n}$ and $\ell_{2}$ and obtained the minimax optimal critical radius. Our work can hence be seen as a generalization to part of Baraud, (2002). In an interesting recent work, Wei and Wainwright, (2020) studied local testing rates on ellipsoids. Here local means that the point of interest under the null is not necessarily the zero point, but can be an arbitrary point in the ellipsoid. Wei and Wainwright, (2020) proved upper and lower bounds based on localized Kolmogorov widths, which are not guaranteed to match in general, but do match for certain examples with Sobolev ellipsoids. In contrast, the present work focuses on more general sets, but the point of interest under the null hypothesis is always anchored at the zero point, and the critical radius is determined based on the global (instead of local) Kolmogorov widths of the set.
@@ 1.3 Organization
This paper is structured as follows. Section 2 introduces the notations and several important concepts used through the paper. Section 3 presents the lower and upper bounds for the signal detection problem and contains the main results of the paper. It also contains the discussion of the phase transitions of the problem. Section 4 is the extension of the main results in Section 3 to the robust $\ell_{p}$ norm testing problem. Section 5 presents the experiment procedures, including the data generation, constraint and adversary setups, the experimental results of the polynomial-time algorithms related to Theorem 3.11 and relevant discussion. Section 6 contains the discussion which compares the minimax rates of robust testing to robust estimation. It also includes some open problems for future works. The pseudo code of the relevant algorithms discussed in the paper and the additional experimental results of Section 5 are provided in Appendix A. We attach necessary probabilistic tools and auxiliary lemmas in Appendix B. The proofs of the lower and upper bounds and other main theorems, lemmas, and corollaries in Section 3, Section 4 and Section 6 are included in Appendix C.
@@ 2.1 Notations and Definitions
To simplify the analysis, we assume that $K\subset\mathcal{X}=\mathbb{R}^{d}$, and therefore $\mu\in\mathbb{R}^{d},X_{i}\in\mathbb{R}^{d}$ through the paper. However, one can extend the results to the case when $K\subset\ell_{2}$ with the same techniques and algorithms with minimal efforts given the assumption $\lim\limits_{k\rightarrow\infty}D_{k}(K)=0$. See Definition 2.2 for further details. We use $\rho$ to denote the value of the $\ell_{2}$ norm of $\mu$ under $H_{1}$. $\left\lVert\cdot\right\rVert_{p}$ means the $\ell_{p}$ norm in $\mathbb{R}^{d}$ for $1\leq p\leq\infty$. When $A$ is a set, $\left|A\right|$ denotes the cardinality of $A$. We use calligraphic font $\mathcal{P},\mathcal{Q},\mathcal{R}\dots$ to represent probability distributions. $\mathcal{P}\ll\mathcal{Q}$ means that $\mathcal{P}$ is absolutely continuous with respect to $\mathcal{Q}$. $\text{TV}(\mathcal{P},\mathcal{Q})$ denotes the total variation distance and $\text{KL}(\mathcal{P},\mathcal{Q})$ for the Kullback-Leibler divergence if $\mathcal{P}\ll\mathcal{Q}$. When the underlying probability distribution is not explicitly specified, we use $\mathbb{P}$ alternatively to denote its probability law in general. We use $\left\langle\cdot,\cdot\right\rangle$ and $(\cdot)^{\top}(\cdot)$ alternatively (possibly with a subscript when needed) to represent the inner product in a Hilbert space. $\text{diag}\{a\}$ represents the diagonal matrix with diagonal entries $a_{i},i=1,2,\dots$. The convenient notation $[n]$ denotes the set $\left\{1,2,\dots,n\right\}$. When $I$ is an index set, $v_{I}$ represents the restriction of $v$ on $I$ if $v$ is a vector and $\mathbf{Y}_{I}$ represents the sub-matrix of $\mathbf{Y}$ with rows selected according to $I$ if $\mathbf{Y}$ is a matrix. $\mathcal{O}$, $\varTheta$, $\varOmega$ notations are used to denote the asymptotic order of rates. The lowercase $c_{1},c_{2},\dots$ denote the universal constants and their specific values may differ from line to line.
@@
Given a set $K\subset\mathbb{R}^{d}$, we say $K$ is a quadratically convex orthosymmetric (QCO) set if it satisfies the following conditions:
 (1), $K$ is convex;
 (2), $K$ is quadratically convex, which means that $K^{2}$ is also convex, where $K^{2}$ is defined as
@@
(3), $K$ is orthosymmetric, which means that if $\theta=\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K$, then $\theta_{\eta}:=\left(\eta_{1}\theta_{1},\dots,\eta_{d}\theta_{d}\right)^{\top}\in K$, where $\eta_{i}\in\left\{-1,1\right\},1\leq i\leq d$.
@@
The definition of QCO set can be naturally extended when $K\subset\ell_{2}$, as we only need to replace $[d]$ by the positive integers $\mathbb{N}^{+}$.
Definition 2.2 (Kolmogorov $k$-width, also known as Kolmogorov $N$-width).
@@
In this work, our analysis will be confined to orthogonal projections. Moreover, though $\mathcal{P}_{k}$ is the set of all orthogonal projection with intrinsic dimension $k$, we discretize this set — only consider the projections that are aligned with the coordinate axes without loss of generality, as shown by the analysis later. For now, let us first assume that the norm leveraged in the definition is the $\ell_{2}$ norm.
Based on the definition, one can derive several basic properties of the Kolmogorov widths. They are non-increasing with respect to $k$, meaning that $D_{k_{1}}(K)\geq D_{k_{2}}(K)$ for any $0\leq k_{1}\leq k_{2}$. In addition, for a finite-dimensional space $\mathcal{X}$ with full dimension $d$, we always have $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}$ and $D_{d}(K)=0$ for any $K\subset\mathcal{X}$. For the more general case where $K\subset\ell_{2}$, we shall assume $\lim\limits_{d\rightarrow\infty}D_{d}(K)=0$. In this paper, we further define $D_{k}(K)=\infty$ when $k<0$ and $D_{k}(K)=0$ for $k>d$ for the convenience of the marginal cases.
The connection between the Kolmogorov widths and our main results is established through the following definition of the first and second dimensions of the optimal projections, which are designed to identify the optimal $k$ satisfying certain marginal conditions determined by the geometric properties of the constraint set $K$ and the corruption process.
@@
$H_{d}(a_{1},\dots,a_{d}):=\left\{x\left\lvert\right.x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d},\left|x_{i}\right|\leq a_{i},a_{i}\geq 0,1\leq i\leq d\right\}.$
Obviously, $H_{d}(a_{1},\dots,a_{d})$ is convex and orthosymmetric. One can verify that $H_{d}^{2}(a_{1},\dots,a_{d})=\left\{x^{2}\left\lvert\right.x\in H_{d}(a_{1},\dots,a_{d})\right\}$ is also a hyperrectangle with each side being $0,a_{i}^{2}],1\leq i\leq d$. Therefore, it is a QCO set. By reindexing if necessary, we may assume that the sequence $\left\{a_{i}\right\}_{i=1}^{d}$ is arranged in decreasing order. The Kolmogorov widths are
@@
Again, we assume $\left\{a_{i}\right\}_{i=1}^{d}$ is arranged in decreasing order, by which the Kolmogorov widths are
@@
The first optimal dimension is $k_{1}^{\star}=\max\limits\left\{j\left\lvert\right.1\leq j\leq d,a_{j}>\frac{\sqrt{j}}{N}\sigma^{2}\right\}$ when $a_{1}>\frac{1}{N}\sigma^{2}$ and 0 otherwise. The second optimal dimension $k_{2}^{\star}=\max\limits\left\{j\left\lvert\right.1\leq j\leq d,a_{j}>\frac{\sqrt{j}\epsilon}{\sqrt{N}}\sigma^{2}\right\}$ when $a_{1}>\frac{\epsilon}{\sqrt{N}}\sigma^{2}$ and 0 otherwise.
@@
When an ellipse is not axes-aligned: (1), the orthosymmetric property is not guaranteed, which can be easily checked; (2), the quadratically convex property is true for $d=2$ and false for $d\geq 3$. To gain the idea of (2), consider an ellipse